maryella47:

Match the following items. 1. (-14) + 81 = 81 + (-14) distributive property 2. • = 1 commutative property of addition 3. 101 + (29 + 417) = (101 + 29) + 417 additive identity 4. (24 + 15) = • 24 + • 15 associative property of addition 5. -72 + 0 = -72 multiplicative inverse

Vocaloid:

there's a bit of info missing from some of these options but I will do my best to provide brief explanations of what these properties are so you can solve this problem distributive: you can distribute a factor across multiple items in a factor to yield an expanded product, ex: A(B+C) = A*B + A*C commutative property of addition: changing the order of terms in addition does not change the sum, ex: A + B = B + A additive identity: adding 0 to a number does not change the number A + 0 = A associative property of addition: changing the parentheses in an addition problem has no effect on the sum (A+B) + C = A + (B + C) multiplicative inverse: multiplying a number by its multiplicative inverse yields 1 as the result (1/A) * A = 1

Vocaloid:

anyway, that's a lot to take in, try to think about each of these explanations and see which of the options is the best match for each property